The bob of the pendulum is 200 g. It is swinging at a height of 1 m. The length of the pendulum is 1 m. Find the values of GPE, KE, and total mechanical energy at a height of 1 m, 0.5 m and 0.25 m and at its lowest point.



1
Expert's answer
2021-11-17T10:37:19-0500

"m = 200g = 0.2kg"

"g = 9.8\frac{m}{s^2}"

"E = GPE +KE"

"GPE = mgh"


"\text{for }h=1m"

"KE = 0"

"GPE = 0.2*9.8*1= 1.96J"

"E = 1.96J"


"\text{for }h=0.5m\newline
E = 1.96J\newline
GPE = 0.2*9.8*0.5=0.98J\newline
KE = E-GPE = 1.96-0.98 =0.98J"


"\text{for }h=0.25m\newline
E = 1.96J\newline
GPE = 0.2*9.8*0.25=0.49J\newline
KE = E-GPE = 1.96-0.49 =1.47J"


"\text{for }h=0m\newline
E = 1.96J\newline
GPE = 0\newline
KE = E-GPE = 1.96J"